OpenStudy (anonymous):

Write the equation of the line that has a slope of -2 and a y-intercept of -6

OpenStudy (anonymous):

y=-2x-6

OpenStudy (anonymous):

Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3

OpenStudy (anonymous):

@jcotter0468

OpenStudy (anonymous):

well, we know its going to look something like this: \[y=-\frac{ 2 }{ 3 }x+b\] but we have to find b. to do that, we can use the point and plug in the x and y values. So we get:\[-2=-\frac{ 2 }{ 3 }(6)+b\] \[-2=-4+b\]\[2=b\] So, the equation is \[y=-\frac{ 2 }{ 3 }x+2\]